myininaya (myininaya):

hey cow this might sound stupid but is there a way to prove a negative times a negative is positive?

myininaya (myininaya):

let put this formally: Let a and b be positive real numbers. Then -a(-b)=ab.

OpenStudy (dumbcow):

assume opposite and prove it false by counterexample i guess: assume (-a)(-b) = -ab counter-example: a=2,b=3 (-2)(-3) = 6 thus premise is false and neg times neg is positive

myininaya (myininaya):

i wish i could think of a proof for this

OpenStudy (dumbcow):

multiplication is just a repeated sum, so the reason for negXneg=pos i think comes from why -(-3) = 3 or 2-(-1)=2+1=3

myininaya (myininaya):

right but i want to prove -(-3)=3

OpenStudy (anonymous):

we know\[-a=ae^{i(2n+1) \pi}\] and \[-b=be^{i(2n+1) \pi}\] so (-a)(-b)=\[ae^{i(2n+1) \pi}be^{i(2n+1) \pi}=abe^{i2(2n+1)\pi}=ab\]

OpenStudy (anonymous):

i think u gao my approach

myininaya (myininaya):

omg seriously wow i think this is good

OpenStudy (dumbcow):

yeah that will work :)

myininaya (myininaya):

:) yes!

OpenStudy (anonymous):

thanks..

myininaya (myininaya):

A>0 B>0 (A)+(-A)=0 B(A)+B(-A)=0 so this proves a negative times a positive is negative since a positvex +negativex=0 (A)+(-A)=0 -B(A)+(-B)(-A)=0 in order for this to be zero one of terms need to be negative and one needs to be positive -B(A) is negative so (-B)(-A)=BA what about this too?

OpenStudy (anonymous):

its also right...but here u considered that +ve*+ve is +ve